# VLG-CBM: Training Concept Bottleneck Models with Vision-Language Guidance

### v2

@@ Abstract
Concept Bottleneck Models (CBMs) provide interpretable prediction by introducing an intermediate Concept Bottleneck Layer (CBL), which encodes human-understandable concepts to explain models’ decision. Recent works proposed to utilize Large Language Models (LLMs) and pre-trained Vision-Language Models (VLMs) to automate the training of CBMs, making it more scalable and automated. However, existing approaches still fall short in two aspects: First, the concepts predicted by CBL often mismatch the input image, raising doubts about the faithfulness of interpretation. Second, it has been shown that concept values encode unintended information: even a set of random concepts could achieve comparable test accuracy to state-of-the-art CBMs. To address these critical limitations, in this work, we propose a novel framework called Vision-Language-Guided Concept Bottleneck Model (VLG-CBM) to enable faithful interpretability with the benefits of boosted performance. Our method leverages off-the-shelf open-domain grounded object detectors to provide visually grounded concept annotation, which largely enhances the faithfulness of concept prediction while further improving the model performance. In addition, we propose a new metric called Number of Effective Concepts (NEC) to control the information leakage and provide better interpretability. Extensive evaluations across five standard benchmarks show that our method, VLG-CBM, outperforms existing methods by at least 4.27% and up to 51.09% on accuracy at NEC=5, and by at least 0.45% and up to 29.78% on average accuracy across different NECs, while preserving both faithfulness and interpretability of the learned concepts as demonstrated in extensive experiments [^1] Our code is available at https://github.com/Trustworthy-ML-Lab/VLG-CBM.
@@ 1 Introduction
    We conduct a series of experiments in Section 5 and demonstrate that our VLG-CBM outperforms existing methods across 5 standard benchmarks by at least 4.27% and up to 51.09% on accuracy at NEC=5, and by at least 0.45% and up to 29.78% on average accuracy across different NECs. Our learned CBM achieves a high sparsity of 0.2% in the final layer even on large datasets including Places365, preserving interpretability even with a large number of concepts. Additionally, we qualitatively demonstrate that our method provides more accurate concept attributions compared to existing methods.
@@ 5.1 Performance comparison
Metrics. As discussed in Section 4, in order to evaluate the final classification power of CBM, the NEC needs to be controlled. Therefore, we measure the following two metrics:
@@
    Accuracy at NEC=5: (Acc@5) This metric is designed to show the performance of CBM which could provide an interpretable prediction. We choose the number 5 so that human users could easily inspect all concepts related to the decision without much effort.
@@
    Average accuracy: To evaluate the trade-off between interpretability and performance, we also calculate the average accuracy under different NECs. In general, higher NEC indicates a more complex model, which may achieve better performance but also hurt interpretability. We choose six different levels: $5,10,15,20,25,30$ and measure the average accuracy.
@@
Metrics Acc@5 Avg. Acc Acc@5 Avg. Acc Acc@5 Avg. Acc Acc@5 Avg. Acc
@@
Metrics Acc@5 Avg. Acc. Acc@5 Avg. Acc. Acc@5 Avg. Acc.
@@
Table 3: Performance comparison with non-CLIP backbone. We compare against LF-CBM[15] and a random baseline, as LM4CV[25], LaBo[27] does not support non-CLIP backbone. The random baseline has 1024 neurons for CIFAR10 and CIFAR100, 512 for CUB, 2048 for Places365, and 4096 for ImageNet.
@@ 6 Conclusion, Potential Limitations and Future work
In this work, we study how to improve the interpretability and performance of concept bottleneck models. We introduce a novel approach VLG-CBM based on both vision and language guidance, which successfully improves both the interpretability and utility of existing CBMs in prior work. Additionally, our theoretical analysis show that information leakage may exist on even in the untrained CBLs, serving the foundations for our proposed new metric (NEC). We show that NEC not only allow fair evaluation of CBMs but also can be used to effectively control information leakage of CBM and ensure interpretability. Extensive experiments on image classification benchmarks demonstrated our VLG-CBM largely outperform previous baselines especially for small NEC, providing more interpretable decisions for users.
@@ Appendix C Ablation Studies
Metrics Acc@5 Avg. Acc. Acc@5 Avg. Acc.
@@
Table C.1: Accuracy at NEC@5 and Average accuracy for different confidence threshold $T$.
@@ C.1 Ablation study for confidence threshold
Confidence threshold $T$ in Eq 2 filters concepts with bounding boxes’ confidence less than $T$. In this experiment, we study the affect of T on the VLG-CBM’s accuracy. The results are shown in Table C.1. We observe that the accuracy at NEM@5 and average accuracy first increases (or stays constant) and then decreases. We attribute this effect to to the fact that as T increases, the number of false-positive decreases leading to better learning of concepts, however, as the number of annotations available for learning a concept decreases.
@@ I.2.1 Ablation on augmentation probability
Crop-to-Concept-Prob Acc@NEC=5 Avg. Acc

### v3

@@ Abstract
Concept Bottleneck Models (CBMs) provide interpretable prediction by introducing an intermediate Concept Bottleneck Layer (CBL), which encodes human-understandable concepts to explain models’ decision. Recent works proposed to utilize Large Language Models and pre-trained Vision-Language Models to automate the training of CBMs, making it more scalable and automated. However, existing approaches still fall short in two aspects: First, the concepts predicted by CBL often mismatch the input image, raising doubts about the faithfulness of interpretation. Second, it has been shown that concept values encode unintended information: even a set of random concepts could achieve comparable test accuracy to state-of-the-art CBMs. To address these critical limitations, in this work, we propose a novel framework called Vision-Language-Guided Concept Bottleneck Model (VLG-CBM) to enable faithful interpretability with the benefits of boosted performance. Our method leverages off-the-shelf open-domain grounded object detectors to provide visually grounded concept annotation, which largely enhances the faithfulness of concept prediction while further improving the model performance. In addition, we propose a new metric called Number of Effective Concepts (NEC) to control the information leakage and provide better interpretability. Extensive evaluations across five standard benchmarks show that our method, VLG-CBM, outperforms existing methods by at least 4.27% and up to 51.09% on Accuracy at NEC=5 (denoted as ANEC-5), and by at least 0.45% and up to 29.78% on average accuracy (denoted as ANEC-avg), while preserving both faithfulness and interpretability of the learned concepts as demonstrated in extensive experiments [^1] Our code is available at https://github.com/Trustworthy-ML-Lab/VLG-CBM.
@@ 1 Introduction
    We conduct a series of experiments in Section 5 and demonstrate that our VLG-CBM outperforms existing methods across 5 standard benchmarks by at least 4.27% and up to 51.09% on Accuracy at NEC=5 (denoted as ANEC-5), and by at least 0.45% and up to 29.78% on average accuracy across different NECs (denoted as ANEC-avg). Our learned CBM achieves a high sparsity of 0.2% in the final layer even on large datasets including Places365, preserving interpretability even with a large number of concepts. Additionally, we qualitatively demonstrate that our method provides more accurate concept attributions compared to existing methods.
@@ 5.1 Performance comparison
Metrics. As discussed in Section 4, in order to evaluate the final classification power of CBM, we should acculate the Accuracy under specified NEC(ANEC). Therefore, we measure the following two metrics:
@@
    Accuracy at NEC=5(ANEC-5): This metric is designed to show the performance of CBM which could provide an interpretable prediction. We choose the number 5 so that human users could easily inspect all concepts related to the decision without much effort.
@@
    Average accuracy(ANEC-avg): To evaluate the trade-off between interpretability and performance, we also calculate the average accuracy under different NECs. In general, higher NEC indicates a more complex model, which may achieve better performance but also hurt interpretability. We choose six different levels: $5,10,15,20,25,30$ and measure the average accuracy.
@@
Metrics ANEC-5 ANEC-avg ANEC-5 ANEC-avg ANEC-5 ANEC-avg ANEC-5 ANEC-avg
@@
Metrics ANEC-5 ANEC-avg ANEC-5 ANEC-avg ANEC-5 ANEC-avg
@@
Table 3: Performance comparison with non-CLIP backbone. We compare against LF-CBM[15] and a random baseline, as LM4CV[25], LaBo[27] does not support non-CLIP backbone. The random baseline has 512 neurons for CUB, 2048 for Places365, and 4096 for ImageNet.
@@ 6 Conclusion, Potential Limitations and Future work
In this work, we study how to improve the interpretability and performance of concept bottleneck models. We introduce a novel approach VLG-CBM based on both vision and language guidance, which successfully improves both the interpretability and utility of existing CBMs in prior work. Additionally, our theoretical analysis show that information leakage may exist on even in the untrained CBLs, serving the foundations for our proposed new NEC-controlled metrics (ANEC-5 and ANEC-avg). We show that NEC not only allow fair evaluation of CBMs but also can be used to effectively control information leakage of CBM and ensure interpretability. Extensive experiments on image classification benchmarks demonstrated our VLG-CBM largely outperform previous baselines especially for small NEC, providing more interpretable decisions for users.
@@ Appendix C Ablation Studies
Metrics ANEC-5 ANEC-avg ANEC-5 ANEC-avg
@@
Table C.1: ANEC-5 and ANEC-avg for different confidence threshold $T$.
@@ C.1 Ablation study for confidence threshold
Confidence threshold $T$ in Eq 2 filters concepts with bounding boxes’ confidence less than $T$. In this experiment, we study the affect of T on the VLG-CBM’s accuracy. The results are shown in Table C.1. We observe that ANEC-5 and ANEC-avg first increases (or stays constant) and then decreases. We attribute this effect to to the fact that as T increases, the number of false-positive decreases leading to better learning of concepts, however, as the number of annotations available for learning a concept decreases.
@@ I.2.1 Ablation on augmentation probability
Crop-to-Concept-Prob ANEC-5 ANEC-avg